find the force of an attraction between two equal but opposite charges of 2*10^-6C if the distance between them is 25cm
1
Expert's answer
2020-07-21T12:56:37-0400

Given points are

Charge is "Q=2*10^{-6} C"

Distance between them is "r=0.25 m"


------------------------------------------------------------------------------

The attraction force between the two charges is

"F=\frac{kQ^2}{r^2}"

"=\frac{(9*10^9)(2*10^{-6})^2}{(0.25)^2}"

"=0.576 N"